# Exploiting Vocabulary Frequency Imbalance in Language Model Pre-training

### v1

@@ Abstract
Large language models are trained with tokenizers, and the resulting token distribution is highly imbalanced: a few words dominate the stream while most occur rarely. Recent practice favors ever-larger vocabularies, but it is unclear where the benefit comes from. To this end, we perform a controlled study that scales the vocabulary of the language model from 2424K to 196196K while holding data, computation, and optimization unchanged. We begin by quantifying the complexity of tokenized text – formalized via Kolmogorov complexity – and show that larger vocabularies reduce this complexity. Above 2424K, every common word is already tokenized as a single token, so enlarging vocabulary only deepens the relative token-frequency imbalance. Word-level loss decomposition shows that larger vocabularies reduce cross-entropy loss almost exclusively by lowering uncertainty on the $2,500$ most frequent words, even though loss on the rare tail rises. Constraining input and output embedding norms to attenuate the impact of token-frequency imbalance reverses the gain, directly demonstrating that the model exploits rather than suffers from imbalance. Because the same frequent words cover roughly $77\%$ of tokens in downstream benchmarks, this training advantage transfers intact. We further show that enlarging model parameters with a fixed vocabulary yields the same frequent-word benefit. Our results recast “bigger vocabularies help” as “lowering complexity of tokenized text helps,” offering a simple, principled knob for tokenizer–model co-design and clarifying the loss dynamics that govern language model scaling in pre-training.
@@ 1 Introduction
A language model incorporates a tokenizer that converts a stream of characters into a series of token IDs, each representing a specific substring (Huang et al., 2025; Rajaraman et al., 2024; Dagan et al., 2024). Tokenization has re-emerged as an powerful tuning knob for language models, with mounting evidence that simply scaling up vocabulary consistently reduces perplexity and improves downstream accuracy across diverse domains and model scales (Tao et al., 2024; Yu et al., 2025). Although this trend is consistently observed in practice, the underlying mechanism responsible for it has yet to be thoroughly investigated.
@@
As vocabulary size grows, adding merge candidates segments frequent words in the training data into a single token and pushes infrequent ones deeper into the long tail, sharpening the relative token-frequency distribution (Schmidt et al., 2025; Liu et al., 2025; Hayase et al., 2024). In unigram models, this simply lowers the Shannon entropy of training data toward its optimal loss (i.e., entropy rate) Rajaraman et al. (2024). This intuition does not transfer to language models as they condition their next-token prediction on the preceding context, which contains a mixture of common and rare tokens Bengio et al. (2000); Brown et al. (2020); OpenAI (2023); Black et al. (2022). Moreover, rare tokens already carry much lower conditional probabilities than marginal ones; mistakes on them incur disproportionately large penalties Pinto et al. (2024); Mircea et al. (2024). Yet rare tokens account for a small share of the entire dataset, making it unclear how a larger vocabulary reallocates capacity between frequent and rare tokens.
@@
In this study, we explore why enlarging vocabulary size improves the performance of language models by expanding vocabulary from 2424K to 196196K. We quantify the complexity of tokenized text using an upper bound on Kolmogorov complexity and first demonstrate that expanding the vocabulary reduces this complexity (§3.3). Above 2424K, frequent words are already encoded as single tokens, so the primary shift is a heightened imbalance in relative token-frequency distribution (§3.4), regardless of dataset quality (§3.5). Furthermore, a word-level loss decomposition shows that a larger vocabulary reduces the loss of frequent words, thereby lowering the model’s overall cross-entropy. (§3.6).
@@
Analytic experiments trace how enlarging the vocabulary changes both training dynamics and generalization behavior through token-frequency imbalance. Our observation suggests that exploiting relative token-frequency imbalance during pre-training is essential for performance (§4.1). In addition, high frequency words in the pre-training corpus largely coincide with those in downstream benchmarks, both in identity and coverage, explaining the close link between training loss and transfer accuracy (§4.2). Scaling the model itself produces the same benefit: it predicts frequent words more accurately, thereby enhancing overall language-model performance (§4.3).
@@
Contribution. We identify that larger vocabularies reduce the complexity of tokenized text, thereby facilitating the model to learn non-i.i.d. patterns in the training data more easily. Our experiments further reveal that beyond a certain size, vocabulary expansion no longer improves segmentation but instead steepens the skewness of the token-frequency distribution. This sharper imbalance alone lowers global cross-entropy by reducing the top $\sim 2{,}500$ frequent words loss despite slight degradation on the rare tail. Through norm-constrained ablations and cross-dataset overlap analyses, we demonstrate that exploiting, rather than mitigating, frequency skew causally reduces cross-entropy and boosts downstream accuracy. Finally, we show that parameter scaling replicates the same benefit as vocabulary scaling, both primarily reduce uncertainty on the same set of frequent tokens.
@@ 2 Motivation
Tokenization – the interface between raw text and the discrete symbols a model actually sees – has resurfaced as a powerful, low-cost lever for improving language model quality (Huang et al., 2025; Rajaraman et al., 2024; Dagan et al., 2024; Hayase et al., 2024). A growing body of evidence finds that simply increasing the size of the tokenizer vocabulary yields systematic gains in perplexity and downstream accuracy across domain and model scales (Tao et al., 2024; Yu et al., 2025). Despite this empirical regularity, the mechanism behind the gain remains underexplored.
The clue may lie in how tokenizers behave as their vocabularies grow. Rajaraman et al. (2024) noted that adding merge candidates segments the most frequent words in the corpus into single tokens, making an i.i.d. model over tokens a closer approximation to inherently non-i.i.d. data. Simultaneously, this pushes rare vocabularies further into the long tail, yielding a markedly more skewed token-frequency distribution (after the usual whitespace pre-tokenization) (Schmidt et al., 2025; Liu et al., 2025). In a unigram model, increasing the relative frequency of common vocabularies reduces Shannon entropy – the theoretical minimum cross-entropy – while a longer tail raises it (Zouhar et al., 2023). Empirically, Rajaraman et al. (2024) confirms that larger vocabularies lower unigram cross-entropy.
Unfortunately, this explanation does not carry over verbatim to neural language models. In language models, every prediction is conditioned on a variable-length context. Errors on rare vocabularies are disproportionately costly because their conditional probabilities are already orders of magnitude below their marginal frequencies Pinto et al. (2024); Mircea et al. (2024), yet they occupy only a tiny fraction of the overall loss. Quantifying how vocabulary growth redistributes modelling capacity between these two regimes, therefore, remains a non-trivial challenge (see the Appendix A for a detailed explanation).
@@ 3 Experiments
    Does enlarging the vocabulary reduce the Kolmogorov complexity of tokenized text (3.3)?
@@ 3.1 Experimental Settings
In these experiments, we train a byte-pair encoding (BPE) tokenizer Sennrich et al. (2016a) and estimate token frequencies using a sample of 10 billion GPT-2 tokens from FineWeb-Edu Penedo et al. (2024) and the entire OpenWebText Gokaslan et al. (2019). For model pre-training, we use approximately 40 billion characters, about 7.5 billion tokens for FineWeb-Edu and 7 billion for OpenWebText with a 49K vocabulary. To compute the metrics below, we also drew an additional 5 billion characters that did not overlap with the training corpus. We report word-level average per-word loss to ensure fair comparison across vocabulary sizes. Whenever a smaller-vocabulary tokenizer splits a word into multiple tokens (i.e., subwords), we sum their individual losses. Our model comprises 85 million non-embedding parameters with pre-layer normalization (pre-LN) Xiong et al. (2020). Training uses AdamW Loshchilov and Hutter (2019) ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$) with a learning rate of $6\times 10^{-4}$ that follows a cosine-decay schedule after a 350 million-token warmup, weight decay of $0.1$, and gradient clipping at $1.0$. Every experiment was repeated with five seeds.
@@ 3.2 Metrics
Kolmogorov complexity $K(X)$ measures the minimal description length of a bitstring $X$ and thus captures its inherent structure and compressibility (Kolmogorov, 1998; Goldblum et al., 2023). However, exact Kolmogorov complexity is uncomputable, since any such procedure would decide the halting problem (Kolmogorov, 1968; Morris et al., 2025; Goldblum et al., 2023). Instead, we calculate a computable upper bound on Kolmogorov complexity, which serves as a practical metric for measuring the complexity of compressed data (Goldblum et al., 2023). As BPE tokenization Sennrich et al. (2016a) is derived from a BPE compression algorithm (Gage, 1994), tokenized text can be viewed as compressed data. For a BPE-tokenized bitstring $X^{N}$, an asymptotic upper bound on $K(X^{N})$ is given by
$K(X^{N})\;\leq\;N\,H(p)\;+\;V\log_{2}N\;+\;O(\log N),$ (1)
where $N$ is the total token count of tokenized text, $H(p)$ is the Shannon entropy of the token distribution, and the $V\log_{2}N$ term accounts for the prefix-free encoding of the token (e.g., token frequency table). Since modern language models are trained on a massive corpus with billions of tokens, the NH(p)N\,H(p) component dominates, yielding the practical approximation $K(X^{N})\approx N\,H(p)$.
@@
$\mathrm{Total\ Loss}(v)\;=\;\sum_{t\in N}\sum_{i=1}^{|t|}\mathbbm{1}(v=t_{i})\,\bigl{[}-\ln p(t_{i}\mid t_{<i})\bigr{]},$ (2)
where $N$ is the set of evaluation documents, $\mathbbm{1}(v=t_{i})$ is the indicator that the $i$th token equals vocabulary $v$, and $\ln p(t_{i}\mid t_{<i})$ is the negative log-likelihood of that token. Total loss measures the sum of negative log-likelihoods for each vocabulary, reflecting its impact on the model’s loss Magnusson et al. (2024).
@@
$\mu(\ell_{v})\;=\;\frac{\mathrm{Total\ Loss}(v)}{T_{v}(N)},$ (3)
where $T_{v}(N)$ is the total count of occurrences of $v$ in the training data. This represents the mean negative log-likelihood across all occurrences of each vocabulary $v$ Magnusson et al. (2024).
@@
$\mathrm{Global\ Cross\ Entropy\ Loss}=\sum_{v}\frac{T_{v}(N)}{T(N)}\,\mu(\ell_{v}),$ (4)
where $T(N)=\sum_{v}T_{v}(N)$ is the total token count in the training data. Global Cross-Entropy Loss reflects the model’s average uncertainty per prediction Jelinek et al. (1977); Magnusson et al. (2024).
Table 1: Upper bound of Kolmogorov complexity ($K(X^{N})$) and NCR for the 45.7 billion-byte FineWeb-Edu corpus with $K(X^{N})$ measured in bytes.
@@
$24\mathrm{K}$ $10.74B$ 0.235
$49\mathrm{K}$ $10.43B$ 0.228
$98\mathrm{K}$ $10.23B$ 0.224
$196\mathrm{K}$ $10.16B$ 0.222
@@ 3.3 Increasing vocabulary size reduces complexity of tokenized text
Naturally occurring data contains a shared structure and inherently low complexity, models can learn effectively from training data, achieving high accuracy across diverse tasks and generalizing to unseen datasets (Goldblum et al., 2023). To quantify how larger vocabularies make text more structured and compressible – better approximating natural data’s low intrinsic complexity – we measure the upper bound of Kolmogorov complexity and the normalized compression ratio (NCR). NCR is a practical compressibility metric $\mathrm{NCR}(x;C)\;=\;\frac{\lvert C(x)\rvert}{\lvert x\rvert}$ where $\lvert x\rvert$ is the byte length of data $x$ and $\lvert C(x)\rvert$ is its size after lossless compression by compressor $C$; in our formulation, $K(X^{N})\approx\lvert C(x)\rvert=N\,H(p)$. Table 1 reports the upper bound on Kolmogorov complexity and NCR for the 45.7 billion-byte FineWeb-Edu corpus. The results show that BPE tokenizers with larger vocabularies yield lower complexity and NCR, resulting in more structured and compressible data. This facilitates models to learn patterns in data more easily, as non-i.i.d. character sequences in text (e.g., words) are segmented as a single token, bringing a tokenized text sequence closer to i.i.d. and simplifying language modeling.
@@ 3.4 Segmentation already saturates; vocabulary growth mainly sharpens frequency skew
In this experiment, relative token-frequency imbalance is quantified by the Jensen–Shannon divergence (JSD) from a uniform distribution of the same vocabulary size for a fair comparison across different vocabulary sizes Zouhar et al. (2023); Dagan et al. (2024); Schmidt et al. (2025). Segmentation efficiency gauges the average token count per word, that is, how many tokens the tokenizer needs, on average, to encode each of the $2,500$ most frequent words. Figure 1(a) shows that JSD grows monotonically with vocabulary size, reflecting an increasingly heavy-tailed relative token-frequency distribution. In contrast, Figure 1(b) indicates that the segmentation efficiency exceeds $95$ % by $24\mathrm{K}$ tokens and plateaus later. In other words, beyond $24\mathrm{K}$, the vocabulary gets larger without providing additional single-token coverage for frequent words. Within the widely used $24\mathrm{K}\!-\!196\mathrm{K}$ range (Schmidt et al., 2024; Ali et al., 2024; Reddy et al., 2025), enlarging the vocabulary chiefly amplifies relative token-frequency imbalance rather than improving segmentation efficiency. This finding tempers the common view that "bigger vocabularies help by approximating word-level tokens" (Rajaraman et al., 2024): that mechanism appears to saturate once frequent words are already tokenized as a single token.

### v2

@@ Abstract
Large language models are trained with tokenizers, and the resulting token distribution is highly imbalanced: a few words dominate the stream while most occur rarely. Recent practice favors ever-larger vocabularies, but it is unclear where the benefit comes from. To this end, we perform a controlled study that scales the vocabulary of the language model from 2424K to 196196K while holding data, computation, and optimization unchanged. We begin by quantifying the complexity of tokenized text – formalized via Kolmogorov complexity – and show that larger vocabularies reduce this complexity. Above 2424K, every common word is already tokenized as a single token, so enlarging vocabulary only deepens the relative token-frequency imbalance. Word-level loss decomposition shows that larger vocabularies reduce cross-entropy loss almost exclusively by lowering uncertainty on the $2,500$ most frequent words, even though loss on the rare tail rises. Same frequent words cover roughly $75\%$ of tokens in downstream benchmarks, this training advantage transfers intact. We further show that enlarging model parameters with a fixed vocabulary yields the same frequent-word benefit. Our results recast “bigger vocabularies help” as “lowering complexity of tokenized text helps,” offering a simple, principled knob for tokenizer–model co-design and clarifying the loss dynamics that govern language model scaling in pre-training.
@@ 1 Introduction
A language model incorporates a tokenizer that converts a stream of characters into a series of token IDs, each representing a specific substring [20, 44, 10]. Tokenization has re-emerged as an powerful tuning knob for language models, with mounting evidence that simply scaling up vocabulary consistently reduces perplexity and improves downstream accuracy across diverse domains and model scales [55, 62]. Although this trend is consistently observed in practice, the underlying mechanism responsible for it has yet to be thoroughly investigated.
@@
As vocabulary size grows, adding merge candidates segments frequent words in the training data into a single token and pushes infrequent ones deeper into the long tail, sharpening the relative token‑frequency distribution [49, 32, 18]. In unigram models, this simply lowers the unigram Shannon entropy of training data toward its optimal loss (i.e., entropy rate) [44]. This intuition does not transfer to language models as they condition their next-token prediction on the preceding context, which contains a mixture of common and rare tokens [2, 7, 39, 6]. Moreover, rare tokens already carry much lower conditional probabilities than marginal ones; mistakes on them incur disproportionately large penalties [42, 36]. Yet rare tokens account for a small share of the entire dataset, making it unclear how a larger vocabulary reallocates capacity between frequent and rare tokens.
@@
In this study, we explore why enlarging vocabulary size improves the performance of language models by expanding vocabulary from 2424K to 196196K. Viewing BPE [51] with its pre-tokenization rules as a lossless compressor [11, 30], we assess how much the tokenized text is compressed against raw text by an upper-bound on Kolmogorov complexity and first illustrate that expanding the vocabulary reduces this complexity (§3.3). Above 2424K, frequent words are already encoded as single tokens, so the primary shift is a heightened imbalance in relative token-frequency distribution (§3.4), regardless of dataset quality (§3.5). Furthermore, a word-level loss decomposition shows that a larger vocabulary reduces the loss of frequent words, thereby lowering the model’s overall cross-entropy (§3.6), regardless of dataset and model size.
@@
Analytic experiments trace how enlarging the vocabulary changes both training dynamics and generalization behavior through token-frequency imbalance. Our observation suggests that high frequency words in the pre-training corpus largely coincide with those in downstream benchmarks, both in identity and coverage, explaining the close link between training loss and transfer accuracy (§4.1). Scaling the model itself produces the same benefit: it predicts frequent words more accurately, thereby enhancing overall language‑model performance (§4.2).
@@
Contribution. We identify that larger vocabularies reduce the complexity of tokenized text, thereby facilitating the model to learn non-i.i.d. patterns in the training data more easily. Our experiments further reveal that beyond a certain size, vocabulary expansion no longer improves segmentation but instead steepens the skewness of the token‐frequency distribution. This sharper imbalance alone lowers global cross‐entropy by reducing the top $\sim 2{,}500$ frequent words loss despite slight degradation on the rare tail. Through cross‐dataset overlap analyses, we demonstrate that exploiting, rather than mitigating, token frequency imbalance causally reduces cross‐entropy and boosts downstream accuracy. Finally, we show that parameter scaling replicates the same benefit as vocabulary scaling, both primarily reduce uncertainty on the same set of frequent tokens.
@@ 2 Motivation
Tokenization – the interface between raw text and the discrete symbols a model actually sees – has resurfaced as a powerful, low-cost lever for improving language model quality [20, 44, 10, 18]. A growing body of evidence finds that simply increasing the size of the tokenizer vocabulary yields systematic gains in perplexity and downstream accuracy across domain and model scales [55, 62]. Despite this empirical regularity, the mechanism behind the gain remains underexplored.
The clue may lie in how tokenizers behave as their vocabularies grow. Rajaraman et al. [44] noted that adding merge candidates segments the most frequent words in the corpus into single tokens, making an i.i.d. model over tokens a closer approximation to inherently non-i.i.d. data. Moreover, they empirically confirm that larger vocabularies lower unigram cross-entropy. Simultaneously, this pushes rare vocabularies further into the long tail, yielding a markedly more skewed token-frequency distribution (after the usual whitespace pre-tokenization) [49, 32].
Unfortunately, this explanation does not carry over verbatim to neural language models. In language models, every prediction is conditioned on a variable-length context. Errors on rare vocabularies are disproportionately costly because their conditional probabilities are already orders of magnitude below their marginal frequencies [42, 36], yet they occupy only a tiny fraction of the overall loss. Quantifying how vocabulary growth redistributes modelling capacity between these two regimes, therefore, remains a non-trivial challenge (see the Appendix A for a detailed explanation).
@@ 3 Experiments
    Does enlarging the vocabulary reduce the Kolmogorov complexity of tokenized text (§3.3)?
@@ 3.1 Experimental Settings
In these experiments, we train a byte-pair encoding (BPE) tokenizer [51] and estimate token frequencies using a sample of 10 billion GPT-2 tokens from FineWeb-Edu [40] and the entire OpenWebText [13]. For model pre-training, we use approximately 40 billion characters, about 7.5 billion tokens for FineWeb-Edu and 7 billion for OpenWebText with a 49K vocabulary. To compute the metrics below, we also drew an additional 5 billion characters that did not overlap with the training corpus. We report word-level average loss to ensure fair comparison across vocabulary sizes. Whenever a smaller-vocabulary tokenizer splits a word into multiple tokens (i.e., subwords), we sum their individual losses. Our model comprises 85 million non-embedding parameters with pre-layer normalization (pre-LN) [59]. Training uses AdamW [33] ($\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$) with a learning rate of $6\times 10^{-4}$ that follows a cosine-decay schedule after a 350 million-token warmup, weight decay of $0.1$, and gradient clipping at $1.0$. Every experiment was repeated with five seeds (See Appendix I).
@@ 3.2 Metrics
Kolmogorov complexity $K(X)$ measures the minimal description length of a bitstring $X$, thus captures its inherent structure and compressibility without assuming any data-generating distribution [27, 14]. This is different from Shannon entropy which is the lower bound of an expected code length for a random variable under its probability distribution (e.g., relative token frequency) [53, 54]. Two concepts are highly correlated each other: shannon entropy approximates an average per-token kolmogorov complexity [16, 37]. We adopt Kolmogorov complexity as a general metric for measuring complexity of tokenized text since different pre-tokenization rules can change token counts dramatically under the same tokenizer settings [32, 49, 45].
Exact Kolmogorov complexity is uncomputable as it reduces to the halting problem [26, 37, 14, 61]. Instead, we calculate a computable upper bound on Kolmogorov complexity, which serves as a practical metric for measuring the complexity of compressed data [14]. As BPE tokenization [51] is derived from a BPE compression algorithm [12], tokenized text can be viewed as compressed data. For a BPE-tokenized bitstring $X^{N}$, an asymptotic upper bound on $K(X^{N})$ is given by
$K(X^{N})\;\leq\;N\,H(p)\;+\;V\log_{2}N\;+\;O(\log N),$
where $N$ is the total token count of tokenized text, $H(p)$ is the unigram Shannon entropy of the token distribution, and the $V\log_{2}N$ term accounts for the prefix-free encoding of the token (e.g., token frequency table). Since each token count is an integer $\leq N$ requiring $\left\lceil\log_{2}\!\left(N+1\right)\right\rceil$ bits, the BPE model is $\leq V\log_{2}N$ bits. The first term upper-bounds the tokenized data size, the second bounds the frequency table size, and the third captures minor logarithmic overhead. Since modern language models are trained on a massive corpus with billions of tokens, the NH(p)N\,H(p) component dominates. We use $K(X^{N})\approx N\,H(p)$, focusing on the complexity of the tokenized sequence itself. Accordingly, we take $H(p)$ as the unigram Shannon entropy, despite it being a loose upper bound, to enable fair token-count comparisons across tokenizers with different vocabulary sizes.
@@
$\mathrm{Total\ Loss}(v)\;=\;\sum_{t\in N}\sum_{i=1}^{|t|}\mathbbm{1}(v=t_{i})\,\bigl[-\ln p(t_{i}\mid t_{<i})\bigr],$
where $N$ is the set of evaluation documents, $\mathbbm{1}(v=t_{i})$ is the indicator that the $i$th token equals vocabulary $v$, and $\ln p(t_{i}\mid t_{<i})$ is the negative log-likelihood of that token. Total loss measures the sum of negative log-likelihoods for each vocabulary, reflecting its impact on the model’s loss [35].
@@
$\mu(\ell_{v})\;=\;\frac{\mathrm{Total\ Loss}(v)}{T_{v}(N)},$
where $T_{v}(N)$ is the total count of occurrences of $v$ in the training data. This represents the mean negative log-likelihood across all occurrences of each vocabulary $v$ [35].
@@
$\mathrm{Global\ Cross\ Entropy\ Loss}=\sum_{v}\frac{T_{v}(N)}{T(N)}\,\mu(\ell_{v}),$
where $T(N)=\sum_{v}T_{v}(N)$ is the total token count in the training data. Global Cross-Entropy Loss reflects the model’s average uncertainty per prediction [23, 35].
Table 1: Upper bound of Kolmogorov complexity ($K(X^{N})$) and NCR for the $45.97$ billion byte FineWeb‑Edu corpus with $K(X^{N})$ measured in bytes.
@@
$24\mathrm{K}$ $10.74B$ 0.234
$49\mathrm{K}$ $10.43B$ 0.227
$98\mathrm{K}$ $10.23B$ 0.223
$196\mathrm{K}$ $10.16B$ 0.221
@@ 3.3 Increasing vocabulary size reduces complexity of tokenized text
Naturally occurring data contains a shared structure and inherently low complexity and models can learn effectively from training data, achieving high accuracy across diverse tasks and generalizing to unseen datasets [14]. To quantify how larger vocabulary size helps the tokenizer approximate natural data’s low intrinsic complexity and produces simpler tokenized sequences, we measure the upper bound of Kolmogorov complexity and the normalized compression ratio (NCR). NCR is a practical compressibility metric $\mathrm{NCR}(x;C)\;=\;\frac{\lvert C(x)\rvert}{\lvert x\rvert}$ where $\lvert x\rvert$ is the byte length of data $x$ and $\lvert C(x)\rvert$ is its size after lossless compression by compressor $C$; in our formulation, $K(X^{N})\approx\lvert C(x)\rvert=N\,H(p)$.
Table 1 reports the upper bound on Kolmogorov complexity and NCR for the $45.97$ billion byte FineWeb‑Edu corpus ($10\mathrm{B}$ GPT-2 tokens). The results show that BPE tokenizers with larger vocabularies yield lower complexity and NCR by segmenting frequent non-i.i.d. character sequences in text (e.g., words) as a single token. This facilitates models to learn non-i.i.d. statistical patterns in natural text more easily by offloading the low-level pattern learning, thereby simplifying language modeling.
@@ 3.4 Segmentation already saturates; vocabulary growth mainly sharpens frequency skew
In this experiment, relative token-frequency imbalance is quantified by the Jensen–Shannon divergence (JSD) from a uniform distribution of the same vocabulary size for a fair comparison across different vocabulary sizes [64, 10, 49]. Segmentation efficiency gauges the average token count per word, that is, how many tokens the tokenizer needs, on average, to encode each of the $2,500$ most frequent words. Figure 1(a) shows that JSD grows monotonically with vocabulary size, reflecting greater token-frequency imbalance. In contrast, Figure 1(b) indicates that the segmentation efficiency exceeds $95$ % by $24\mathrm{K}$ tokens and plateaus later. In other words, beyond $24\mathrm{K}$, the vocabulary gets larger without providing additional single-token coverage for frequent words. Within the widely used $24\mathrm{K}\!-\!196\mathrm{K}$ range [48, 1, 46], enlarging the vocabulary chiefly amplifies relative token-frequency imbalance rather than improving segmentation efficiency. This finding tempers the common view that "bigger vocabularies help by approximating word-level tokens" [44]: that mechanism appears to saturate once frequent words are already tokenized as a single token.